OpenStudy (anonymous):

The sum of an integer and its reciprocal is 5/2. Find the integer.

OpenStudy (anonymous):

x is our integer

OpenStudy (anonymous):

1/x is its reciprocal x + 1/x = 5/2 Do you know the next step?

OpenStudy (anonymous):

x+1/x=5/2 you need to write 5/2 in the form of x+1/x. 5/2 can be written as a mixed fraction. This fraction is 2+1/2. You have now written 5/2 in the form of x+1/x, and in your answer x was equal to 2. Therefore, 2 is your answer!